OpenStudy (anonymous):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x). A. 7 + 4x B. 4 – x C. –x – 4 D. –4 – 7x

OpenStudy (aaronq):

f(x) + g(x)= 3x – 2 + 6 – 4x =3x-4x-2+6 can you finish it?

OpenStudy (anonymous):

Add 4x to both sides?

OpenStudy (aaronq):

no no, now that the "like" terms are grouped, you simply add or subtract them. 3x-4x=-x -2+6=4 =-x+4 which is also 4-x

OpenStudy (anonymous):

Thank you(:

OpenStudy (aaronq):

no problem !